if cos 20 - sin 20=p, find the value of cos 40 in terms of p

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

DEar Beverly

cos 20 - sin 20=p

square it

cos220 + sin220 -2sin20cos20 =p2

sin40 =1-p2

and cos40 =(1-sin240)1/2

                   =(2p2-p4)1/2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin